a continuous random variable T has the following probability density function.

f(t) = { 32t2/128, -4≤t<4

0, otherwise


a. find E(t)

b. find Var(T)

c. Var(2T+3)



1
Expert's answer
2021-06-06T16:19:27-0400

Solution:

"f(t) = \{ \dfrac{32t^2}{128}, -4≤t<4
\\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0, otherwise"

(a):

"E(t)=\int_{-4}^4t(\dfrac{32t^2}{128})=\int_{-4}^4(\dfrac{t^3}{4})
\\=[\dfrac{t^4}{16}]_{-4}^4
\\=\dfrac{4^4}{16}-\dfrac{(-4)^4}{16}
\\=0"

(b):

"E(t^2)=\int_{-4}^4(\dfrac{t^4}{4})
\\=[\dfrac{t^5}{20}]_{-4}^4
\\=\dfrac{4^5}{20}-\dfrac{(-4)^5}{20}
\\=2(\dfrac{4^5}{20})=102.4"

Now, "Var[T]=\sigma^2=E(X^2)-[E(X)]^2=102.4-0^2=102.4"

(c):

"Var[2T+3]=4\ Var[T]=4(102.4)=409.6"